The energy of a single photon is 2.038x10-18 J.  What is the energy of one mole of these photons, in kilojoules?
Answer #1

Energy of one mole = Energy of one photon*Avogadro no.

= ( 2.038×10-18)(6.022×1023)

( Avogadro no.=6.022×1023 /mol)

= 12.27×105 j/mol

. = 12.27×105×10-3  (1j=10-3 kj)

Energy of one mole = 12.27×102 kj/mol

  Energy of one mole = 1227 kj/mol